OpenStudy (anonymous):

Find the product of (x + 3)^2.

OpenStudy (anonymous):

x^2 + 3^2

OpenStudy (anonymous):

x^2 + 9

OpenStudy (mr.math):

You can do it two ways: \[1) (x+3)(x+3)=x^2+3x+3x+9=x^2+6x+9\].

OpenStudy (anonymous):

(x + 3)^2 (x + 3)(x + 3) x^2 + 3x +3x + 9 x^2 + 6x + 9

OpenStudy (mr.math):

Or apply the formula \((a+b)^2=a^2+2ab+b^2\).

OpenStudy (pokemon23):

Bob06 finally...